OpenStudy (anonymous):

Find the x-coordinates where f '(x) = 0 for f(x) = 2x + sin(4x) in the interval [0, π].

OpenStudy (solomonzelman):

You need to find f'(x), and then to set f'(x)=0, to solve for x-values at which f'(x)=0.

OpenStudy (solomonzelman):

What is the derivative of\(\tiny~\)?? f(x) = 2x + sin(4x)

OpenStudy (anonymous):

Sorry had to do some chores

OpenStudy (anonymous):

4cos(4x)+2 this is the derivative

OpenStudy (anonymous):

cos(4x) = -1/2 4x=arc cos -1/2 I get to here and then confused where to go next

OpenStudy (solomonzelman):

f'(x)=4cos(4x)+2 0=4cos(4x)+2 4cos(4x)=-2 cos(4x)=-1/2

OpenStudy (solomonzelman):

you can just solve this graphically https://www.desmos.com/calculator/up3hfxrzld

OpenStudy (anonymous):

Yeah I know that but I feel my teacher will want work shown

OpenStudy (anonymous):

Actually Ill just do the intersect in my calculator and explain the logic

OpenStudy (anonymous):

would it be possible for you to explain this as well http://prntscr.com/9rxh67 I do not understand it a bit.